# Multiscale Modeling of Vacancy-Cluster Interactions and Solute Clustering Kinetics in Al-Mg-Zn Alloys

### v1

@@ Results
Figure 7: Comparison between CD and KMC simulations under infinitely fast (instantaneous) cooling conditions. (a) Time evolution of system energy per atom. Results are shown for KMC (blue dashed, light blue shade represents the standard error from three individual KMC runs), CD without vacancy-cluster interaction corrections (NVC, green), and CD with vacancy-cluster interaction corrections (VC, orange). (b)-(d) The distributions of cluster concentrations $c^{\text{clu}}$ as a function of cluster size $N^{\text{clu}}$ at the time of (b) 1 s, (c) 30 s, and (d) 880 s after the starting of the simulations.
@@
To validate the predictive capability of the CD model, particularly its incorporation of vacancy-cluster interaction corrections in solute diffusivity, we directly compare the CD results with KMC simulations under infinitely fast (instantaneous) cooling conditions. Fig. 7a presents the time evolution of the system energy per atom, showing three datasets: KMC simulations, CD without vacancy-cluster interaction corrections (CD-NVC), and CD with corrections (CD-VC). The CD-VC model includes the effects of dynamic residual vacancy concentration, $c^{\text{vac,res}}_{\text{dyn}}$, which captures vacancy supersaturation and its influence on solute diffusivity $D^{\text{X}}$. Since solute diffusivity directly governs condensation and evaporation rates in the CD framework, accurately incorporating $c^{\text{vac,res}}_{\text{dyn}}$ allows CD-VC simulations to closely reproduce the energy evolution obtained from KMC over time. In addition, Figs. 7b-d compare the distributions of cluster concentrations, $c^{\text{clu}}$, as a function of cluster size, $N^{\text{clu}}$, obtained from CD-VC predictions and KMC simulations at different time points: 1 s, 30 s, and 880 s. These comparisons demonstrate strong agreement between the CD-VC and KMC models in predicting the temporal evolution of cluster size distributions across a wide range of times and temperatures.
@@
Finally, the combination of KMC simulations and the CD model, informed by the calculated dynamic vacancy concentrations $c^{\text{vac,res}}_{\text{dyn}}$ (which directly influence solute diffusivity), is employed to simulate the long-time natural aging process following various cooling conditions, beginning from the as-quenched state. In this hybrid approach, KMC simulations are first used to model the quenching process until the temperature reaches approximately 300 K (corresponding to 10 seconds in our cases). At this point, the system energy evolution plateaus as shown in Fig. 1, indicating that most vacancies in the supercell are effectively trapped by solute clusters, and the system reaches a metastable state. Subsequently, the CD model is applied to simulate the long-term behavior of solute cluster evolution during natural aging. Fig. 8a presents the evolution of residual vacancy concentrations, $c^{\text{vac,res}}_{\text{dyn}}$, during natural aging at 300 K for both fast and slow cooling scenarios. These scenarios correspond to the two temperature profiles shown in Fig. 6a. The initial value of $c^{\text{vac,res}}_{\text{dyn}}$ at the onset of aging (i.e., at time $t=10$ s, marking the end of quenching) strongly depends on the cooling rate. Faster cooling limits the time available for solute clusters to form and trap vacancies, resulting in higher initial values of $c^{\text{vac,res}}_{\text{dyn}}$, as shown in Fig. 6b.
@@
During the subsequent natural aging process, the dynamic residual vacancy concentration, $c^{\text{vac,res}}_{\text{dyn}}$, continues to decrease over time for both cooling scenarios. This decline reflects the gradual annihilation of excess vacancies at various sinks, particularly growing solute clusters. Although dislocations and GB are also known to act as vacancy sinks, they typically become significant only at much longer timescales (on the order of $10^{4}$-$10^{5}$ s after quenching [4]) and are therefore not explicitly included in the present model. The time-dependent evolution of $c^{\text{vac,res}}_{\text{dyn}}$ plays a critical role in determining solute diffusivity and thus strongly influences the kinetics of solute clustering during natural aging. Figs 8b and 8c show the cluster concentration distributions, $c^{\text{clu}}$, as a function of cluster size, $N^{\text{clu}}$, predicted by the CD simulations at various natural aging times: as-quenched (AQ), 1 hour, 6 hours, and 12 hours after the quenching. These results are shown for both fast and slow cooling conditions. Under the fast cooling condition (Fig. 8b), $c^{\text{clu}}(N^{\text{clu}})$ distributions corresponding to small quenched-in clusters are initially present. These clusters continue to grow and coarsen during natural aging due to the elevated diffusivity enabled by higher vacancy supersaturation in the Al matrix. In contrast, the slow cooling condition (Fig. 8c) leads to the formation of larger initial clusters and a substantially reduced concentration of mobile vacancies. As a result, the cluster size distribution evolves much more slowly during natural aging. These results demonstrate that the rate of cluster evolution during aging is strongly influenced by the preceding thermal history, with faster cooling leading to more rapid diffusion and clustering kinetics due to a greater degree of vacancy supersaturation.
Discussions
In this study, we developed an integrated computational framework that combines atomistic KMC simulations with continuum-scale CD modeling, centered on an absorbing Markov chain model, to quantitatively investigate vacancy-cluster interactions in Al-Mg-Zn alloys and their impact on long-time natural aging kinetics. While the absorbing Markov chain approach has been previously applied to address energy basin challenges in computational materials science [19, 20, 21], we specifically adapted it here to explore vacancy trapping in multicomponent alloy systems. Our model accurately predicts vacancy escape times ($t_{\text{esc}}$) from Mg-Zn-rich clusters across a wide range of cluster sizes and configurations. It is parameterized using energetics from surrogate models trained on first-principles calculations (Figs. 2b-c) [54, 30], explicitly incorporating both thermodynamic driving forces via local site binding energies ($E_{\text{bind}}^{\text{site}}$) and kinetic barriers through local bond migration energies ($E_{\text{mig}}^{\text{bond}}$). This approach provides a comprehensive description of the so-called “vacancy prison” effect induced by solute clusters (Fig. 2e). Importantly, we rigorously validated the model against direct KMC simulations (Fig. 4), demonstrating its accuracy and the robustness of the simplified PEL assumption, even in the presence of atomic rearrangements within multicomponent clusters.
@@
The absorbing Markov chain model developed in this work represents a significant advancement in understanding vacancy-cluster interactions, particularly in reconciling previously conflicting descriptions of effective vacancy binding energies. Although the importance of vacancy-cluster binding in solute clustering has been widely recognized, earlier studies [33, 55, 56, 57, 58, 59, 10, 60, 61, 8, 9, 53, 11] have provided contrasting models for how the binding energy evolves with cluster size. For instance, Zurob and Seyedrezai [33] proposed a linear relationship between vacancy binding energy and cluster size, suggesting $E^{\text{clu}}_{\text{bind}}\propto N^{\text{clu}}E^{\text{X}}_{\text{bind}}$, where $N^{\text{clu}}$ is the number of solute atoms in the cluster and $E^{\text{X}}_{\text{bind}}$ is the vacancy binding energy of an individual solute atom. In contrast, Soisson et al. [10] and Vincent et al. [9] studied Cu clusters in $\alpha$-Fe and found that the effective binding energy saturates to a value determined by the difference in vacancy formation energies between the matrix and the precipitate: $E^{\text{clu}}_{\text{bind}}=E_{\text{form}}^{\text{vac,ppt}}-E_{\text{form}}^{\text{vac,mx}}$.
@@
The Markov chain model developed in this study plays a crucial role in enabling mesoscale simulations of early-stage precipitation kinetics. A key challenge in accurately modeling this behavior lies in quantifying the dynamic concentration of vacancies in the presence of solute clusters formed during quenching. This is addressed by incorporating the effective vacancy binding energies (Eq. 13) into the CD framework (Fig. 6). Our predicted vacancy concentrations show good agreement with experimental observations, which consistently report that faster cooling rates lead to higher residual vacancy concentrations [62, 51]. This vacancy concentration correction is essential, as it directly governs solute diffusivity ($D^{\text{X}}$), nucleation rate ($dN/dt$), and growth kinetics ($dr/dt$) during aging [4, 5, 36]. The predictive capability of this approach is validated by excellent agreement with direct KMC simulations (Fig. 7). Furthermore, long-time CD simulations (Fig. 8) reveal that quenching conditions have a lasting impact on the rate of cluster evolution during natural aging. This result provides a quantitative explanation for extensive experimental findings that slow cooling rates degrade mechanical properties such as strength and toughness [63, 64, 65, 66, 67, 68]. The degradation arises because slow cooling promotes the formation of coarse, quench-induced precipitates, which consume both solute atoms and excess vacancies, limiting the formation of fine, strengthening precipitates during subsequent aging [69, 62, 51].
@@
Despite the advances enabled by our integrated multiscale framework, some limitations remain in the current implementation. In particular, our CD simulations, although informed by KMC and Markov chain model results, are performed using a fixed Zn/Mg ratio of 2. This simplification is adopted to avoid the computationally prohibitive task of exploring the full two-dimensional compositional space defined by the Mg and Zn atomic fractions ($n_{\text{Mg}}$, $n_{\text{Zn}}$) in cluster formation enthalpy calculations. The selected ratio is physically meaningful, corresponding to the minimum cluster formation enthalpy pathway (see Supplementary Note 2) and consistent with recent thermodynamic assessments of Mg-Zn clustering in Al alloys [70]. While this approach does not capture the dynamic evolution of the Zn/Mg ratio, which is known to vary with alloy composition and thermomechanical processing conditions [71, 72, 73, 74], it remains a reasonable approximation within the scope of this study. Specifically, our focus is on vacancy trapping effects and their impact on long-term aging kinetics, where this simplification is justified by two key observations. First, at the atomistic level, we find that vacancy trapping behavior is comparable between clusters derived from SA simulations with a fixed Zn/Mg ratio of 2 and those from KMC simulations with varying Zn/Mg ratios (see Supplementary Note 3). Second, at the mesoscale, our simplified CD model accurately reproduces the system energy evolution and cluster size distributions predicted by direct KMC simulations (Fig. 7). These results collectively support the validity of the fixed-ratio approximation for the purposes of this study.
@@
Additionally, while our CD simulations effectively model long-term cluster evolution kinetics, they remain inherently limited by assumptions from classical nucleation theory (CNT) and simplified growth models, such as single-atom addition mechanisms. These simplifications restrict the ability of CD to capture complex phenomena such as coarsening, cluster morphology evolution, and microstructural interactions. To overcome these limitations, larger-scale simulation techniques, such as PF modeling [22, 23, 24], are essential. PF methods are well-suited for simulating microstructural evolution and morphological changes across larger spatial and temporal scales. The predictive capability of PF simulations, however, depends critically on the accurate parameterization of the free energy functional and mobility terms. Our integrated computational framework provides key inputs for such models: surrogate models trained on first-principles data yield reliable thermodynamic quantities for free energy landscapes, while the dynamic vacancy concentrations derived from our approach offer corrections to solute diffusivities and mobilities required in PF formulations. This integration opens pathways for more comprehensive, large-scale simulations of precipitation and microstructure evolution in engineering alloys.
@@ Methods
The details of the KMC simulation methods are provided in our recent studies and are briefly summarized below [30]. The most common approach to KMC simulations is the first-order residence time algorithm [75]. In this algorithm, we consider $z=12$ first nearest neighbors of a vacancy. Assuming each neighboring atom vibrates at a typical crystal frequency ($\nu_{i}=1\times 10^{13}$ Hz), it results in a transition probability per unit time $r_{i}=\nu_{i}\exp(-E^{\text{bond}}_{\text{mig},i}/k_{B}T)$, with $E^{\text{bond}}_{\text{mig},i}$ representing the migration energy barrier. However, low-barrier events near the solute clusters often lead to forward and backward vacancy jumps, causing frequent ’flicker’ events that considerably slow down first-order KMC algorithms, particularly at low temperatures. We employed a second-order residence time algorithm [38, 39, 41], which bypasses flicker events by considering two-step vacancy jumps. For solute clustering studies, KMC simulations were performed in a $30\times 30\times 30$ FCC supercell of Al-2.86 at.%Mg-2.38 at.%Zn-0.001 at.%vacancy, representative of 7075 Al alloy compositions. The solute concentrations are within the range of compositions of the 7075 Al alloys. Initial configurations were equilibrated at 800 K (typical solutionizing temperature for Al alloys) via CMC for approximately 109 steps. During simulated quenching, temperatures were updated at each timestep according to predefined cooling profiles. For the absorbing Markov chain model verification, simulations employed the same $17\times 17\times 17$ FCC supercell containing clusters obtained from SA and one vacancy, conducted at a higher temperature of 450 K for simulation efficiency.
@@
Figure 9: The transition graph of an absorbing Markov chain with 5 states, where states 1, 2, and 3 are transient states and states 4 and 5 are absorbing states.
A Markov chain is a model that describes a system transitioning between a set of states with certain probabilities. It assumes that the next state depends only on the current state, not the past sequence of states. The system is defined by a set of positions (all possible states), a transition matrix (probabilities of moving from one state to another). Monte Carlo (MC) simulations closely relate to Markov chains as both model stochastic system evolution via states with Markovian transition probabilities. The transition matrix $\boldsymbol{P}$ is a matrix of all transition rates $P_{ij}$. Here $P_{ij}$ is the probability for state $i$ to state $j$. The summation of all elements in each row should be 1, we have $\sum_{j}P_{ij}=1$.
An absorbing Markov chain is a type of Markov chain that includes at least one absorbing state, which is defined as a state that, once entered, cannot be exited. This means the probability of remaining in the absorbing state is 1, i.e., $P_{ii}=1$, and the probability of transitioning to any other state is 0, i.e., $P_{ij}=0$ for $i\neq j$. In contrast, a transient state is a state that can be exited, making it possible to leave and potentially never return. The transition matrix $\boldsymbol{P}$ of an absorbing Markov chain can be expressed in canonical form as:
@@
Fig. 9 shows an example graph of an absorbing Markov chain with 5 states, where states 1, 2, and 3 are transition states and states 4 and 5 are absorbing states. Eq. 14 for this specific graph is:
$\boldsymbol{P}_{\text{example}}=\begin{pmatrix}0&P_{12}&P_{13}&P_{14}&0\\ P_{21}&0&P_{23}&0&P_{25}\\ P_{31}&P_{32}&0&P_{34}&P_{35}\\ 0&0&0&P_{44}&0\\ 0&0&0&0&P_{55}\end{pmatrix}.$ (15)
@@
CD simulations bridge atomistic simulations and macroscopic approaches by tracking the evolution of clusters in an alloy [44, 45, 46, 47, 48]. In this formalism, clusters with size of $N^{\text{clu}}=n$ atoms were explicitly tracked via their concentration $c_{n}$, governed by a set of coupled ordinary differential equations known as the master equations:
@@
where $\alpha_{n}$ represents the evaporation rate (the rate of cluster shrinkage from size $n$ to $n-1$) and $\beta_{n}$ denotes the condensation rate (the rate of cluster growth from size $n$ to $n+1$). The concentration of monomers (single atoms), $c_{1}$, is constrained by the conservative condition of the total solute concentration $c_{\text{tot}}$:
@@
Although the accurate description of multicomponent clusters in Al-Mg-Zn alloys requires treating $\Delta G_{n}$ as a function of both $n_{\text{Mg}}$ and $n_{\text{Zn}}$, we adopt a fixed Zn/Mg ratio of 2 to reduce complexity and ensure computational efficiency. This Zn/Mg ratio in Mg-Zn clusters is almost consistent with the recent findings of Liu et al. [70]. The enthalpy of formation $\Delta H_{n}$ is modeled using a fitting function [78]:
@@
where $A=-0.969$ eV, $B=-0.250$ eV, $C=0.605$ eV, $D=-0.044$ eV, $p=0.952$, and $q=63$ are fitting parameters from the SA simulation results. For the entropy term $\Delta S_{n}$, we only consider the configurational entropy contributions when solute atoms are in dilute solid solutions:
$\Delta S_{n}\approx k_{B}\left[\frac{n}{3}\ln\left(c_{\text{tot}}^{\text{Mg}}\right)+\frac{2n}{3}\ln\left(c_{\text{tot}}^{\text{Zn}}\right)\right]$ (21)
$\beta_{n}$ were explicitly modeled as [79, 80]:
@@
where angle $\theta$ denotes the nucleation and growth direction in the $(n_{\text{Mg}},n_{\text{Zn}})$ space. With Zn/Mg ratio of 2, $\theta=\arctan 2$. $\boldsymbol{B}$ is the matrix characterizing the diffusion-limited condensation process for the critical cluster [45]:
@@
Wu et al. [81] provided first-principles solute diffusion constants, $D^{\text{X}}_{0}$ for Mg of $9.4182\times 10^{-2}$ cm2/s and Zn of $1.0614\times 10^{-1}$ cm2/s, and solute diffusion activation energy, $Q^{X}$, for Mg and Zn in Al matrix of $1.2467$ eV and $1.1812$ eV, respectively. $c^{\text{vac,res}}_{\text{dyn}}$ is calculated based on Eq. 13 using parameters $\tilde{E}_{\text{bind}}^{\text{clu}}$ obtained from KMC and absorbing Markov chain models. $c^{\text{vac,Al}}_{\text{eq}}$ is the equilibrium vacancy concentration:
$c^{\text{vac,Al}}_{\text{eq}}=c_{0}\exp(\frac{S^{\text{vac,Al}}}{k_{B}})\exp(-\frac{H^{\text{vac,Al}}}{k_{B}T})$ (25)

### v2

@@ Results
Figure 7: Comparison between CD and KMC simulations under infinitely fast (instantaneous) cooling conditions. (a) Time evolution of system energy per atom. Results are shown for KMC (blue dashed line with light blue shades that represent the standard errors from three individual KMC simulations), CD without vacancy-cluster interaction corrections (CD-NVC, green solid line), and CD with vacancy-cluster interaction corrections (CD-VC, orange solid line). (b)-(d) The distributions of cluster concentrations $c^{\text{clu}}$ as a function of cluster size $N^{\text{clu}}$ at the time of (b) 1 s, (c) 30 s, and (d) 880 s after the starting of the simulations predicted by both KMC (blue dashed lines) and CD-VC (orange solid lines) method.
@@
To validate the predictive capability of the CD model, particularly its incorporation of vacancy-cluster interaction corrections in solute diffusivity, we directly compare the CD results with KMC simulations under infinitely fast (instantaneous) cooling conditions. Fig. 7a presents the time evolution of the system energy per atom, showing three datasets: KMC simulations, CD without vacancy-cluster interaction corrections (CD-NVC), and CD with corrections (CD-VC). The CD-VC model includes the effects of the dynamic residual vacancy concentration, $c^{\text{vac,res}}_{\text{dyn}}$, which captures vacancy supersaturation and its influence on solute diffusivity $D^{\text{X}}$. Since solute diffusivity directly governs the condensation and evaporation rates in the CD framework, accurately incorporating $c^{\text{vac,res}}_{\text{dyn}}$ allows CD-VC simulations to closely reproduce the energy evolution obtained from KMC over time. In addition, Figs. 7b-d compare the distributions of cluster concentrations, $c^{\text{clu}}$, as a function of cluster size, $N^{\text{clu}}$, obtained from CD-VC predictions and KMC simulations at different time points: 1 s, 30 s, and 880 s after the starting of the simulations. These comparisons demonstrate strong agreement between the CD-VC and KMC models in predicting the temporal evolution of cluster size distributions across a wide range of times and temperatures.
@@
Finally, the combination of KMC simulations and the CD model, informed by the calculated dynamic vacancy concentrations $c^{\text{vac,res}}_{\text{dyn}}$, is employed to simulate the long-time natural aging process following various cooling conditions, beginning from the as-quenched state. In this hybrid approach, KMC simulations are first used to model the quenching process until the temperature reaches approximately 300 K (corresponding to 10 seconds after the starting of the simulations in our cases). At this point, the system energy evolution plateaus as shown in Fig. 1, indicating that vacancies in the supercell are effectively trapped by solute clusters, and the system reaches a metastable state. Subsequently, the CD model is applied to simulate the long-term behavior of solute cluster evolution during natural aging.
Fig. 8a presents the evolution of residual vacancy concentrations in the Al matrix, $c^{\text{vac,res}}_{\text{dyn}}$, during natural aging at 300 K for both fast and slow cooling scenarios. These scenarios correspond to the two temperature profiles shown in Fig. 6a. The initial value of $c^{\text{vac,res}}_{\text{dyn}}$ at the onset of aging (i.e., at time $t=10$ s, marking the end of quenching) strongly depends on the cooling rate. Faster cooling limits the time available for solute clusters to form and trap vacancies, resulting in higher initial values of $c^{\text{vac,res}}_{\text{dyn}}$, which are consistent with the values of $c^{\text{vac,res}}_{\text{dyn}}$ shown in Fig. 6b for these two cooling cases.
@@
During the subsequent natural aging process, $c^{\text{vac,res}}_{\text{dyn}}$ continues to decrease over time for both cooling scenarios. This decline reflects the gradual annihilation of excess vacancies at various sinks, particularly growing solute clusters. Although dislocations and GB are also known to act as vacancy sinks, they typically become significant only at much longer timescales (on the order of $10^{4}$-$10^{5}$ s after quenching [4]) and are therefore not explicitly included in the present model. The time-dependent evolution of $c^{\text{vac,res}}_{\text{dyn}}$ plays a critical role in determining solute diffusivity and thus strongly influences the kinetics of solute clustering during natural aging. Figs 8b and 8c show the cluster concentration distributions, $c^{\text{clu}}$, as a function of cluster size, $N^{\text{clu}}$, predicted by the CD simulations at various natural aging times: as-quenched (AQ), 1 hour, 6 hours, and 12 hours after the quenching. These results are shown for both fast and slow cooling conditions as shown in Fig. 6a. Under the fast cooling condition (Fig. 8b), $c^{\text{clu}}(N^{\text{clu}})$ distributions corresponding to small quenched-in clusters are initially present. These clusters continue to grow and coarsen during natural aging due to the elevated diffusivity enabled by higher vacancy supersaturation in the Al matrix. In contrast, the slow cooling condition (Fig. 8c) leads to the formation of larger initial as-quenched clusters and a substantially reduced concentration of mobile vacancies. As a result, the cluster size distribution evolves much more slowly during natural aging. These results demonstrate that the rate of cluster evolution during aging is strongly influenced by the preceding thermal history, with faster cooling leading to more rapid diffusion and clustering kinetics due to a greater degree of vacancy supersaturation.
Discussion
@@
The absorbing Markov chain model developed in this work represents a significant advancement in understanding vacancy-cluster interactions, particularly in reconciling previously conflicting descriptions of effective vacancy binding energies. As mentioned in the Introduction, earlier studies [38, 63, 64, 65, 66, 67, 15, 68, 69, 13, 14, 62, 16] have provided contrasting models for how the binding energy evolves with cluster size. For instance, Zurob and Seyedrezai [38] proposed a linear relationship between vacancy binding energy and cluster size, suggesting $E^{\text{clu}}_{\text{bind}}\propto N^{\text{clu}}E^{\text{X}}_{\text{bind}}$, where $N^{\text{clu}}$ is the number of solute atoms in the cluster and $E^{\text{X}}_{\text{bind}}$ is the vacancy binding energy of an individual solute atom. In contrast, Soisson et al. [15] and Vincent et al. [14] studied Cu clusters in $\alpha$-Fe and found that the effective binding energy saturates to a value determined by the difference in vacancy formation energies between the matrix and the precipitate: $E^{\text{clu}}_{\text{bind}}=E_{\text{form}}^{\text{vac,ppt}}-E_{\text{form}}^{\text{vac,mx}}$.
@@
The Markov chain model developed in this study plays a crucial role in enabling mesoscale simulations of early-stage precipitation kinetics. A key challenge in accurately modeling this behavior lies in quantifying the dynamic concentration of vacancies in the presence of solute clusters formed during quenching. This is addressed by incorporating the effective vacancy binding energies (Eq. 13) into the CD framework (Fig. 6). Our predicted vacancy concentrations show good agreement with experimental observations, which consistently report that faster cooling rates lead to higher residual vacancy concentrations [70, 60]. This vacancy concentration correction is essential, as it directly governs solute diffusivity, nucleation rate, and growth kinetics during aging [4, 5, 41]. The predictive capability of this approach is validated by excellent agreement with direct KMC simulations (Fig. 7). Furthermore, long-time CD simulations (Fig. 8) reveal that quenching conditions have a lasting impact on the rate of cluster evolution during natural aging. This result provides a quantitative explanation for extensive experimental findings that slow cooling rates degrade mechanical properties such as strength and toughness [71, 72, 73, 74, 75, 76]. The degradation arises because slow cooling promotes the formation of coarse, quench-induced precipitates, which consume both solute atoms and excess vacancies, limiting the formation of fine, strengthening precipitates during subsequent aging [77, 70, 60].
@@
Despite the advances enabled by our integrated multiscale framework, some limitations remain in the current implementation. In particular, our CD simulations, although informed by KMC and Markov chain model results, are performed using a fixed Zn/Mg ratio of 2. This choice dose not reflect inherent limitations of CD method, as well-developed theories for modeling clustering in non-stoichiometric compounds already exist for decades [78, 79, 80, 81, 82, 51]. Instead, this ratio was adopted to avoid the computationally prohibitive task of exploring the full two-dimensional compositional space defined by the number of Mg and Zn atoms ($n_{\text{Mg}}$, $n_{\text{Zn}}$) in cluster formation enthalpy calculations. The selected ratio is physically meaningful, corresponding to the minimum cluster formation enthalpy pathway (see Supplementary Note 2) and consistent with the Zn/Mg ratio observed in large GP-I zones after prolonged aging [59], as well as with recent thermodynamic evaluations of Mg-Zn clustering in Al alloys [83]. While this approach does not capture the dynamic evolution of the Zn/Mg ratio, which is known to vary with alloy composition and thermomechanical processing conditions [84, 85, 86], it remains a reasonable approximation within the scope of this study. Specifically, our focus is on vacancy trapping effects and their impact on long-term aging kinetics, where this simplification is justified by two key observations. First, at the atomistic level, we find that vacancy trapping behavior is comparable between clusters derived from SA simulations with a fixed Zn/Mg ratio of 2 and those from KMC simulations with varying Zn/Mg ratios (see Supplementary Note 3). Second, at the mesoscale, our simplified CD model accurately reproduces the system energy evolution and cluster size distributions predicted by direct KMC simulations (Fig. 7). These results collectively support the validity of the fixed-ratio approximation for the purposes of this study.
@@
Additionally, while our CD simulations effectively model long-term cluster evolution kinetics, they remain inherently limited by assumptions from classical nucleation theory (CNT) and simplified growth models, such as single-atom addition mechanisms. These simplifications restrict the ability of CD to capture complex phenomena such as coarsening, cluster morphology evolution, and microstructural interactions. To overcome these limitations, larger-scale simulation techniques, such as phase field (PF) modeling [27, 28, 29], are essential. PF methods are well-suited for simulating microstructural evolution and morphological changes across larger spatial and temporal scales. The predictive capability of PF simulations, however, depends critically on the accurate parameterization of the free energy functional and mobility terms. Our integrated computational framework provides key inputs for such models: surrogate models trained on first-principles data yield reliable thermodynamic quantities for free energy landscapes, while the dynamic vacancy concentrations derived from our approach offer corrections to solute diffusivities and mobilities required in PF formulations. This integration opens pathways for more comprehensive, large-scale simulations of precipitation and microstructure evolution in engineering alloys.
@@ Methods
The details of the KMC simulation methods are provided in our recent studies and are briefly summarized below [35]. The most common approach to KMC simulations is the first-order residence time algorithm [87]. In this algorithm, we consider $z=12$ first nearest neighbors of a vacancy. Assuming each neighboring atom vibrates at a typical crystal frequency ($\nu_{i}=1\times 10^{13}$ Hz), it results in a transition probability per unit time $r_{i}=\nu_{i}\exp(-E^{\text{bond}}_{\text{mig},i}/k_{B}T)$, with $E^{\text{bond}}_{\text{mig},i}$ representing the migration energy barrier. However, low-barrier events near the solute clusters often lead to forward and backward vacancy jumps, causing frequent ’flicker’ events that considerably slow down first-order KMC algorithms, particularly at low temperatures. We employed a second-order residence time algorithm [43, 44, 46], which bypasses flicker events by considering two-step vacancy jumps. For solute clustering studies, KMC simulations were performed in a $30\times 30\times 30$ FCC supercell of Al-2.86 at.%Mg-2.38 at.%Zn (one vacancy in a whole simulation supercell), representative of 7075 Al alloy compositions. The solute concentrations are within the range of compositions of the 7075 Al alloys. The initial configurations were equilibrated at 800 K (typical solutionizing temperature for Al alloys) via CMC for approximately 109 steps. During simulated quenching, temperatures were updated at each timestep according to predefined cooling profiles. For the absorbing Markov chain model verification, simulations employed the same $17\times 17\times 17$ FCC supercell containing clusters obtained from SA simulations and one vacancy, conducted at a higher temperature of 450 K for simulation efficiency.
@@
Figure 9: A simplified example of the transition graph of an absorbing Markov chain with 5 states, where states 1, 2, and 3 are transient states and states 4 and 5 are absorbing states.
A Markov chain is a model that describes a system transitioning between a set of states with certain probabilities. It assumes that the next state depends only on the current state, not the past sequence of states. Mathematically, it is fully defined by a set of possible states and a transition matrix, whose elements represent the probabilities of moving from a given current state to each possible next state. The transition matrix $\boldsymbol{P}$ is a matrix of all transition rates $P_{ij}$, which is the probability for state $i$ to state $j$. The summation of all elements in each row should be 1, we have $\sum_{j}P_{ij}=1$.
An absorbing Markov chain is a type of Markov chain that includes at least one absorbing state, which is defined as a state that, once entered, cannot be exited. This means the probability of an absorbing state remaining in the absorbing state is 1, i.e., $P_{ii}=1$, and the probability of this absorbing state transitioning to any other state is 0, i.e., $P_{ij}=0$ for $i\neq j$. In contrast, a transient state is a state that can be exited, making it possible to leave and potentially never return. The transition matrix $\boldsymbol{P}$ of an absorbing Markov chain can be expressed in canonical form as:
@@
Fig. 9 shows a simplified example graph of an absorbing Markov chain with 5 states, where states 1, 2, and 3 are transition states and states 4 and 5 are absorbing states. Eq. 14 for this specific graph is:
$\boldsymbol{P}_{\text{example}}=\begin{pmatrix}0&P_{12}&P_{13}&P_{14}&0\\ P_{21}&0&P_{23}&0&P_{25}\\ P_{31}&P_{32}&0&P_{34}&P_{35}\\ 0&0&0&P_{44}&0\\ 0&0&0&0&P_{55}\end{pmatrix}$, (15)
where $P_{44}=P_{55}=1$ because of the absorbing states.
@@
The CD simulations bridge atomistic simulations and macroscopic approaches by tracking the evolution of clusters in an alloy [50, 51, 52, 53, 54]. In this formalism, clusters with a size of $N^{\text{clu}}=n$ atoms were explicitly tracked via their concentration $c_{n}$, governed by a set of coupled ordinary differential equations known as the master equations:
@@
where $\alpha_{n}$ represents the evaporation rate (the rate of cluster shrinkage from size $n$ to n1n-1) and $\beta_{n}$ denotes the condensation rate (the rate of cluster growth from size $n$ to $n+1$). The concentration of monomers (single atoms), $c_{1}$, is constrained by the conservative condition of the total solute concentration $c_{\text{tot}}$:
@@
Although the accurate description of multicomponent clusters in Al-Mg-Zn alloys requires treating $\Delta G_{n}$ as a function of both $n_{\text{Mg}}$ and $n_{\text{Zn}}$, to reduce complexity and ensure computational efficiency, we adopt a fixed Zn/Mg ratio of 2, with the rationale explained in the Discussion section. The enthalpy of formation $\Delta H_{n}$ is modeled using a fitting function [90]:
@@
where A=0.969A=-0.969 eV, B=0.250B=-0.250 eV, $C=0.605$ eV, D=0.044D=-0.044 eV, $p=0.952$, and $q=63$ are fitting parameters from the SA simulation results. For the entropy term $\Delta S_{n}$, we only consider the configurational entropy contributions when solute atoms are in dilute solid solutions:
$\Delta S_{n}\approx k_{B}\left[\frac{n}{3}\ln\left(c_{\text{tot}}^{\text{Mg}}\right)+\frac{2n}{3}\ln\left(c_{\text{tot}}^{\text{Zn}}\right)\right],$ (21)
where $c_{\text{tot}}^{\text{Mg}}$ and $c_{\text{tot}}^{\text{Zn}}$ are concentrations of each solute element present in the entire alloy, regardless of whether they are distributed as isolated solute atoms or within clusters. The condensation rate $\beta_{n}$ was explicitly modeled as [79, 81]:
@@
where angle $\theta$ denotes the nucleation and growth direction in the $(n_{\text{Mg}},n_{\text{Zn}})$ space. With Zn/Mg ratio of 2, $\theta=\arctan 2$. Here $B_{ij}$ in Eq. 22 is the element of matrix $\boldsymbol{B}$, which that characterizes the diffusion-limited condensation process for the critical cluster [51]:
@@
Wu et al. [91] provided first-principles solute diffusion constants, $D^{\text{X}}_{0}$ for Mg of $9.4182\times 10^{-2}$ cm2/s and Zn of $1.0614\times 10^{-1}$ cm2/s, and solute diffusion activation energy (including both vacancy formation and migration energies), $Q^{X}$, for Mg and Zn in Al matrix of $1.2467$ eV and $1.1812$ eV, respectively. $c^{\text{vac,res}}_{\text{dyn}}$ is calculated based on Eq. 13 using parameters $\tilde{E}_{\text{bind}}^{\text{clu}}$ obtained from KMC and absorbing Markov chain models. $c^{\text{vac,Al}}_{\text{eq}}$ is the equilibrium vacancy concentration:
$c^{\text{vac,Al}}_{\text{eq}}=c_{0}\exp\left(\frac{S^{\text{vac,Al}}}{k_{B}}\right)\exp\left(-\frac{H^{\text{vac,Al}}}{k_{B}T}\right)$ (25)